Venus Ltd made a takeover bid for all the issued voting shares of Mars Ltd offering 3 Venus Ltd shares for every 2 Mars Ltd shares. The offer from Venus Ltd was accepted by 90% of the shareholders of Mars Ltd. At acquisition date, 1 January 20X7, the following information was available. Mars Ltd at 31 December 20X6 $’000 Share Capital (200 000 shares) 200 Capital Profits Reserve 300 Retained Profits 150 Total Equity 650 Market value of each company's shares at 1 January 20X7 Venus Ltd $3.20 Mars Ltd $4.00 In Mars Ltd's accounting records, land was stated at $160 000 below its cost to the economic entity (i.e. fair value at acquisition date). This was accounted for as a consolidation adjustment. The rate of company income tax was 30%. What was the goodwill on acquisition date? A) $206 000 B) $394 000 C) $178 000 D) $135 000
Answer #1

The correct answer is D as per below:

As on 1st Jan 20*7. The calculation of net assets are

Share capital 200

Capital Reserve 300

Retained Profits 150

Total Equity 650

Land Cost 160

Net Assets 810

Purchase consideration is 200000*90/100 = 180000*(3/2)* 3.2= $ 864000 or $ 864

90% of Net Assets of Mars = 810*90% = $ 729

Goodwill is 864-729= $ 135 or $ 135000